# Optimal Estimation of Temperature

### v4

@@ IV Estimation of temperature in repeated sampling
where $\sigma_{M}(E)\equiv\sum_{\bm{i}}\delta\left(E-E_{\bm{i}}\right)d_{\bm{i}}$ denotes the density of states of the sampling ensemble. Because Eq. (18) also belongs to the exponential family, Eq. (19) is the UMVUE of $\beta^{k}$ in repeated sampling for system without upper energy bound, according to the RBLS theorem [63, 70, 71]. For system with upper energy bound, the result is similar to that obtained earlier. Especially, for $k=\pm 1$, we have
@@
For the sampling ensemble, the $M$-dependence of density of states, entropy, and temperature implies that sample size should be treated as a independent state variable. In the large-$M$ limit, the sampling ensemble corresponds to Hill’s nanothermodynamics with $\mathcal{N}$-copies, as realized through the replica trick [41, 42, 43, 44, 45, 21]. Accordingly, thermodynamic quantities of entropy and temperature coincide with their counterparts in nanothermodynamics.
@@
$\begin{split}\sqrt{M}\Delta\hat{\beta}_{M}\Delta E&\geq\sqrt{M}\Delta\hat{\beta}_{B,M}\Delta E\\ &=1+\frac{1}{4M}(\mu_{3})^{2}+O\left((MN)^{-2}\right).\end{split}$ (25)
Similarly, because $\hat{T}_{G,M}$ is the UMVUE of $T$ in $M$ samplings, and we obtain
@@
$\begin{split}\frac{\sqrt{M}\Delta\hat{T}_{M}\Delta E}{T^{2}}&\geq\frac{\sqrt{M}\Delta\hat{T}_{G,M}\Delta E}{T^{2}}\\ &=1+\frac{T^{2}}{4MC^{3}}\left(\frac{\text{d}C}{\text{d}T}\right)^{2}+O\left((MN)^{-2}\right)\end{split}$ (27)
in the large $M$ limit. As $M\to\infty$, the Gaussian-type temperature fluctuation and the Cramér-Rao bound are recovered due to CLT. In Table 1, we also list the relevant quantities in the estimation of temperature for the single sampling and repeated sampling.
@@
In FIG.5, we show the probability distributions of the estimators $\hat{T}_{G,M}$ and $\hat{T}_{B,M}=(\partial\ln[\sigma_{M}(E)\epsilon]/\partial E)^{-1}$ in finite-sized system for different $M$. For small $M$, they both follow a non-Gaussian distribution. The non-Gaussian probability distribution of $\hat{T}_{G,M}$ exactly corresponds to the steady state described in Ref. [49], which is a quasi-equilibrium state [50, 52, 53]. As $M$ increases, a Gaussian distribution naturally emerges. Meanwhile, regardless of the value of $M$, the optimal estimation of temperature is always unbiased ( $\left\langle\hat{T}_{G,M}\right\rangle=1$, while $\left\langle\hat{T}_{B,M}\right\rangle\neq 1$). Moreover, for $M=10$, the three estimators $\hat{T}_{G,1},\hat{T}_{G,5},\hat{T}_{G,10}$ are all unbiased. And the variance of $\hat{T}_{G,10}$ are minimum (its distribution is narrowest). Hence, our work provides a foundation for further experimental studies on the non-Gaussian distribution of temperature and the thermodynamic relation between entropy and temperature in finite-sized systems.
@@ V Conclusion and Outlook
Our work provides a theoretical foundation for further exploring the use of many-body systems to construct high-precision thermometers. The achievable bound for the ETU, the non-Gaussian distribution of temperature and the thermodynamic relation between entropy and temperature enable experimental testing in finite-sized systems, such as neutral atom arrays [74, 75] and biochemical oscillators [76, 77]. For thermometry, the optimal estimation of temperature may improve the validity of the measurement and the sampling efficiency in many-body systems.
Acknowledgements.
This work was supported by the Innovation Program for Quantum Science and Technology (Grant No. 2024ZD0301000), the Science Challenge Project (No. TZ2025017), the National Natural Science Foundation of China (Grant No. 12405046), and the Science Foundation of Zhejiang Sci-Tech University (Grants No. 23062088-Y and No. 23062181-Y).
@@ Appendix A Unbiased estimator of $\beta^{k}$
We prove that $\hat{\beta}^{(k)}$ is an unbiased estimator of $\beta^{k}$. Let us consider the case $k>0$. For $k=1$, $\text{d}_{E_{i}}\sigma(E_{i})=\text{d}\sigma(E_{i})/\text{d}E_{i}$, we have [78]
$\begin{split}\langle\hat{\beta}^{(1)}\rangle&=\frac{1}{Z}\sum_{i\in I}\frac{1}{\sigma(E_{i})}\frac{\text{d}\sigma(E_{i})}{\text{d}E_{i}}d_{i}\text{e}^{-\beta E_{i}}\\ &=\frac{1}{Z}\sum_{i\in I}\frac{1}{\sigma(E_{i})}\frac{\text{d}\sigma(E_{i})}{\text{d}E_{i}}d_{i}\text{e}^{-\beta E_{i}}\int_{-\infty}^{\infty}\text{d}E\delta(E-E_{i})\\ &=\frac{1}{Z}\int_{-\infty}^{\infty}\text{d}E\text{e}^{-\beta E}\frac{1}{\sigma(E)}\frac{\text{d}\sigma(E)}{\text{d}E}\sum_{i\in I}\delta(E-E_{i})d_{i}\\ &=\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}{\text{e}^{-\beta E}}\text{d}\sigma(E)\\ &=\left.\frac{\text{e}^{-\beta E}\sigma(E)}{Z}\right|^{\infty}_{E_{\text{min}}}-\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}\sigma(E)\text{d}\text{e}^{-\beta E}\\ &=\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}\sigma(E)\beta\text{e}^{-\beta E}\text{d}E\\ &=\beta,\end{split}$ (28)
@@
$\begin{split}\langle\hat{\beta}^{(n+1)}\rangle&=\frac{1}{Z}\sum_{i\in I}\frac{1}{\sigma(E_{i})}\frac{\text{d}\left[\text{d}^{(n)}_{E_{i}}\sigma(E_{i})\right]}{\text{d}E_{i}}d_{i}\text{e}^{-\beta E_{i}}\\ &=\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}\text{e}^{-\beta E}\text{d}\left[\text{d}^{n}_{E}\sigma(E)\right]\\ &=\left.\frac{\text{e}^{-\beta E}\text{d}^{n}_{E}\sigma(E)}{Z}\right|^{\infty}_{E_{\text{min}}}-\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}\left[\text{d}^{n}_{E}\sigma(E)\right]\text{d}\text{e}^{-\beta E}\\ &=\beta\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}\left[\text{d}^{n}_{E}\sigma(E)\right]\text{e}^{-\beta E}\text{d}E\\ &=\beta^{n+1},\end{split}$ (29)
@@
Next, let us consider the case $k<0$. For k=1k=-1, $\text{d}_{E_{i}}^{-1}\sigma(E_{i})=\int_{E_{\text{min}}}^{E_{i}}\text{d}E^{\prime}\sigma(E^{\prime})$, we have
$\begin{split}\langle\hat{\beta}^{(-1)}\rangle=&\frac{1}{Z}\sum_{i\in I}\frac{\text{d}_{E_{i}}^{-1}\sigma(E_{i})}{\sigma(E_{i})}d_{i}\text{e}^{-\beta E_{i}}\\ =&\frac{-1}{Z\beta}\int_{E_{\text{min}}}^{\infty}\left(\text{d}_{E}^{-1}\sigma(E)\right)\text{d}\text{e}^{-\beta E}\\ =&\left.\frac{-\text{d}_{E}^{-1}\sigma(E)\text{e}^{-\beta E}}{Z\beta}\right|^{\infty}_{E_{\text{min}}}\\ &+\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}\frac{1}{\beta}\text{e}^{-\beta E}\text{d}\left(\text{d}_{E}^{-1}\sigma(E)\right)\\ =&\frac{1}{\beta}\int_{E_{\text{min}}}^{\infty}\sigma(E)\frac{\text{e}^{-\beta E}}{Z}\text{d}E\\ =&\beta^{-1}.\end{split}$ (30)
@@ Appendix B The density of states of an analogous low-temperature ideal Fermi gas
$E=\frac{3}{5}N\varepsilon_{F}+\frac{\pi^{2}}{4}N\varepsilon_{F}\left(\frac{T}{\varepsilon_{F}}\right)^{2}=E_{\text{min}}+aNT^{2},$ (31)
@@
$S=\frac{\pi^{2}TN}{2\varepsilon_{F}}=bNT,$ (32)
where $\varepsilon_{F}$ is generally referred to as the Fermi energy of the system, $a=\pi^{2}/4\varepsilon_{F}$ and $b=\pi^{2}/2\varepsilon_{F}$. Then, $T$ is obtained from Eq. (31):
$T=\sqrt{\frac{E-E_{\text{min}}}{aN}}.$ (33)
Using the Boltzmann relation $S=\ln\sigma(E)$ and Eq. (33), the density of states $\sigma(E)$ can be expressed as:
$\sigma(E)=\text{e}^{S}=\text{e}^{bNT}=\text{e}^{\frac{b}{\sqrt{a}}\sqrt{N(E-E_{\text{min}})}}.$ (34)
For simplicity, we set $b/\sqrt{a}=1$ when plotting FIG. 1. However, the qualitative results are unaffected regardless of the value of $b/\sqrt{a}$. The above approach provides an approximation for the density of states of a low-temperature ideal Fermi gas in the limit of $N\to\infty$. For finite $N$, a more detailed discussion of the density of states can be found in Refs. [80, 81].
@@ Appendix C Relative biases of $\beta$ and $T$ in a two-level system
Firstly, for systems with finite $E_{\text{max}}$, Eq. (29) is modified to
$\begin{split}\langle\hat{\beta}^{(1)}\rangle&=\left.\frac{\sigma(E)\text{e}^{-\beta E}}{Z}\right|^{E_{\text{max}}}_{E_{\text{min}}}-\frac{1}{Z}\int_{E_{\text{min}}}^{E_{\text{max}}}\sigma(E)\text{d}\text{e}^{-\beta E}\\ &=\frac{\sigma(E_{\text{max}})\text{e}^{-\beta E_{\text{max}}}-\sigma(E_{\text{min}})\text{e}^{-\beta E_{\text{min}}}}{Z}+\beta.\end{split}$ (35)
@@
$\sigma(E)=\frac{\Gamma(N+1)}{\varepsilon\Gamma\left(\frac{N}{2}-\frac{E}{\varepsilon}+1\right)\Gamma\left(\frac{N}{2}+\frac{E}{\varepsilon}+1\right)}$ (36)
and $Z=(\text{e}^{\beta\varepsilon/2}+\text{e}^{-\beta\varepsilon/2})^{N}$. Then, Eq. (35) is
$\langle\hat{\beta}^{(1)}\rangle=\frac{\text{e}^{\frac{-\beta N\varepsilon}{2}}-\text{e}^{\frac{\beta N\varepsilon}{2}}}{\varepsilon\left(\text{e}^{\frac{\beta\varepsilon}{2}}+\text{e}^{\frac{-\beta\varepsilon}{2}}\right)^{N}}+\beta,$ (37)
@@
$\begin{split}\Phi_{\beta}&=\left|\frac{\langle\hat{\beta}^{(1)}\rangle-\beta}{\beta}\right|\\ &=\left|\frac{\text{e}^{\frac{-\beta N\varepsilon}{2}}-\text{e}^{\frac{\beta N\varepsilon}{2}}}{\beta\varepsilon\left(\text{e}^{\frac{\beta\varepsilon}{2}}+\text{e}^{\frac{-\beta\varepsilon}{2}}\right)^{N}}\right|\\ &\approx\frac{1}{\beta\varepsilon(1+\text{e}^{-|\beta|\varepsilon})^{N}}\\ &=\frac{1}{\beta\varepsilon}\text{e}^{-N(\ln(1+\text{e}^{-|\beta|\varepsilon}))}\\ &\sim\text{e}^{-\alpha_{1}N},\end{split}$ (38)
where the approximation used applies $\text{e}^{-N|\beta|\varepsilon}\ll 1$ with large $N$. In FIG. 6a, $\alpha_{1}$ decreases exponentially with increasing inverse temperature. For a particular $N$, $\Phi_{\beta}$ is non-negligible, which indicates that the unbiased estimator $\hat{\beta}^{(1)}$ remains unbiased except at extremely low temperatures, as given in Eq. (6).
Figure 6: Decay rate varies as NN when $E_{\text{max}}$ is finite. Here, the density of states of an $N$ non-interacting two-level system is Eq. (36) and we set the spacing of energy levels $\varepsilon=1$.
Secondly, for systems with finite $E_{\text{max}}$, Eq. (30) is modified to
$\begin{split}\langle\hat{\beta}^{(-1)}\rangle=&\left.\frac{-\text{d}_{E}^{-1}\sigma(E)\text{e}^{-\beta E}}{Z\beta}\right|^{E_{\text{max}}}_{E_{\text{min}}}\\ &+\frac{1}{Z}\int_{E_{\text{min}}}^{E_{\text{max}}}\frac{1}{\beta}\text{e}^{-\beta E}\text{d}\left(\text{d}_{E}^{-1}\sigma(E)\right)\\ =&\frac{\Omega(E_{\text{max}})\text{e}^{-\beta E_{\text{max}}}-\Omega(E_{\text{min}})\text{e}^{-\beta E_{\text{min}}}}{\beta Z}+\beta^{-1}.\end{split}$ (39)
In an $N$ non-interacting two-level system, we have $\Omega(E_{\text{min}})=\int_{E_{\text{min}}}^{E_{\text{min}}}\sigma(E)\text{d}E=0$ and $\Omega(E_{\text{max}})=\int_{E_{\text{min}}}^{E_{\text{max}}}\sigma(E)\text{d}E=Z|_{\beta=0}$. Then, with the help of Eq. (36), Eq. (39) is
$\langle\hat{\beta}^{(-1)}\rangle=\frac{2^{N}\text{e}^{\frac{-\beta N\varepsilon}{2}}}{\beta\left(\text{e}^{\frac{\beta\varepsilon}{2}}+\text{e}^{\frac{-\beta\varepsilon}{2}}\right)^{N}}+\beta^{-1},$ (40)

### v5

@@ 4 Estimation of temperature in repeated sampling
where $\sigma_{M}(E)\equiv\sum_{\bm{i}}\delta\left(E-E_{\bm{i}}\right)d_{\bm{i}}$ denotes the density of states of the sampling ensemble. And Eq. (19) is the UMVUE of $\beta^{k}$ in repeated sampling for system without upper energy bound, according to the RBLS theorem [54, 61, 62]. For system with upper energy bound, the result is similar to that obtained earlier. Especially, for $k=\pm 1$, we have
@@
For the sampling ensemble, the $M$-dependence of density of states, entropy, and temperature implies that sample size should be treated as a independent state variable. In the large-$M$ limit, the sampling ensemble corresponds to Hill’s nanothermodynamics with $M$-copies, as realized through the replica trick [19, 20, 21, 22, 23]. Accordingly, thermodynamic quantities of entropy and temperature coincide with their counterparts in nanothermodynamics.
@@
$\begin{split}\sqrt{M}\Delta\hat{\beta}_{M}\Delta E&\geq\sqrt{M}\Delta\hat{\beta}_{B,M}\Delta E=1+\frac{1}{4M}(\mu_{3})^{2}+O\left((MN)^{-2}\right).\end{split}$ (25)
Similarly, we obtain
@@
$\begin{split}\frac{\sqrt{M}\Delta\hat{T}_{M}\Delta E}{T^{2}}&\geq\frac{\sqrt{M}\Delta\hat{T}_{G,M}\Delta E}{T^{2}}=1+\frac{T^{2}}{4MC^{3}}\left(\frac{\text{d}C}{\text{d}T}\right)^{2}+O\left((MN)^{-2}\right)\end{split}$ (27)
in the large $M$ limit. As $M\to\infty$, the Gaussian-type temperature fluctuation and the Cramér-Rao bound are recovered due to the CLT. In Table 1, we also list the relevant quantities in the estimation of temperature for the single sampling and repeated sampling.
@@
In figure 5, we show the probability distributions of the estimators $\hat{T}_{G,M}$ and $\hat{T}_{B,M}=(\partial\ln[\sigma_{M}(E)\epsilon]/\partial E)^{-1}$ in finite-sized system for different $M$. For small $M$, they both follow a non-Gaussian distribution. The non-Gaussian probability distribution of $\hat{T}_{G,M}$ exactly corresponds to the quasi-equilibrium state in Ref. [38] (also see [40, 41, 42]), which can be regarded as a realization of superstatistics [32, 33, 34] in finite-sized systems. As $M$ increases, a Gaussian distribution naturally emerges. Meanwhile, regardless of the value of $M$, the optimal estimation of temperature is always unbiased ( $\left\langle\hat{T}_{G,M}\right\rangle=1$, while $\left\langle\hat{T}_{B,M}\right\rangle\neq 1$). And for $M=10$, the variance of $\hat{T}_{G,10}$ are minimum among the set of $\{\hat{T}_{G,1},\hat{T}_{G,5},\hat{T}_{G,10}\}$. Hence, our work provides a foundation for further experimental studies on the non-Gaussian distribution of temperature and the thermodynamic relation between entropy and temperature in finite-sized systems.
@@ 5 Conclusion and Outlook
The achievable bound for the ETU, the non-Gaussian distribution of temperature and the thermodynamic relation between entropy and temperature enable experimental testing in finite-sized systems, such as neutral atom arrays [65, 66] and biochemical oscillators [67, 68]. For thermometry, the optimal estimation of temperature may improve the validity of the measurement and the sampling efficiency in many-body systems with a finite number of measurements..
[…]
This work was supported by the Science Challenge Project (No. TZ2025017), the Quantum Science and Technology-National Science and Technology Major Project (Grant No.2024ZD 0301000) , the National Natural Science Foundation of China (Grant No. 12405046), and the Science Foundation of Zhejiang Sci-Tech University (Grants No. 23062088-Y and No. 23062181-Y).
[…]
All data that support the findings of this study are included within the article [76].
[…]
@@ Appendix A Unbiased estimator of $\beta^{k}$
We prove that $\hat{\beta}^{(k)}$ is an unbiased estimator of $\beta^{k}$. Let us consider the case $k>0$. For $k=1$, where $\text{d}_{E_{i}}\sigma(E_{i})=\text{d}\sigma(E_{i})/\text{d}E_{i}$, we have [69]
$\begin{split}\langle\hat{\beta}^{(1)}\rangle&=\frac{1}{Z}\sum_{i\in D}\frac{1}{\sigma(E_{i})}\frac{\text{d}\sigma(E_{i})}{\text{d}E_{i}}d_{i}\text{e}^{-\beta E_{i}}\\ &=\frac{1}{Z}\sum_{i\in D}\frac{1}{\sigma(E_{i})}\frac{\text{d}\sigma(E_{i})}{\text{d}E_{i}}d_{i}\text{e}^{-\beta E_{i}}\int_{-\infty}^{\infty}\text{d}E\delta(E-E_{i})\\ &=\frac{1}{Z}\int_{-\infty}^{\infty}\text{d}E\text{e}^{-\beta E}\frac{1}{\sigma(E)}\frac{\text{d}\sigma(E)}{\text{d}E}\sum_{i\in D}\delta(E-E_{i})d_{i}\\ &=\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}{\text{e}^{-\beta E}}\text{d}\sigma(E)\\ &=\left.\frac{\text{e}^{-\beta E}\sigma(E)}{Z}\right|^{\infty}_{E_{\text{min}}}-\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}\sigma(E)\text{d}\text{e}^{-\beta E}\\ &=\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}\sigma(E)\beta\text{e}^{-\beta E}\text{d}E\\ &=\beta,\end{split}$ (A1)
@@
$\begin{split}\langle\hat{\beta}^{(n+1)}\rangle&=\frac{1}{Z}\sum_{i\in D}\frac{1}{\sigma(E_{i})}\frac{\text{d}\left[\text{d}^{(n)}_{E_{i}}\sigma(E_{i})\right]}{\text{d}E_{i}}d_{i}\text{e}^{-\beta E_{i}}\\ &=\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}\text{e}^{-\beta E}\text{d}\left[\text{d}^{n}_{E}\sigma(E)\right]\\ &=\left.\frac{\text{e}^{-\beta E}\text{d}^{n}_{E}\sigma(E)}{Z}\right|^{\infty}_{E_{\text{min}}}-\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}\left[\text{d}^{n}_{E}\sigma(E)\right]\text{d}\text{e}^{-\beta E}\\ &=\beta\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}\left[\text{d}^{n}_{E}\sigma(E)\right]\text{e}^{-\beta E}\text{d}E\\ &=\beta^{n+1},\end{split}$ (A2)
@@
Next, let us consider the case $k<0$. For k=1k=-1, where $\text{d}_{E_{i}}^{-1}\sigma(E_{i})=\int_{E_{\text{min}}}^{E_{i}}\text{d}E^{\prime}\sigma(E^{\prime})$, we have
$\begin{split}\langle\hat{\beta}^{(-1)}\rangle=&\frac{1}{Z}\sum_{i\in D}\frac{\text{d}_{E_{i}}^{-1}\sigma(E_{i})}{\sigma(E_{i})}d_{i}\text{e}^{-\beta E_{i}}\\ =&\frac{-1}{Z\beta}\int_{E_{\text{min}}}^{\infty}\left(\text{d}_{E}^{-1}\sigma(E)\right)\text{d}\text{e}^{-\beta E}\\ =&\left.\frac{-\text{d}_{E}^{-1}\sigma(E)\text{e}^{-\beta E}}{Z\beta}\right|^{\infty}_{E_{\text{min}}}+\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}\frac{1}{\beta}\text{e}^{-\beta E}\text{d}\left(\text{d}_{E}^{-1}\sigma(E)\right)\\ =&\frac{1}{\beta}\int_{E_{\text{min}}}^{\infty}\sigma(E)\frac{\text{e}^{-\beta E}}{Z}\text{d}E\\ =&\beta^{-1}.\end{split}$ (A3)
@@ Appendix B The density of states of an analogous low-temperature ideal Fermi gas
$E=\frac{3}{5}N\varepsilon_{F}+\frac{\pi^{2}}{4}N\varepsilon_{F}\left(\frac{T}{\varepsilon_{F}}\right)^{2}=E_{\text{min}}+aNT^{2},$ (B1)
@@
$S=\frac{\pi^{2}TN}{2\varepsilon_{F}}=bNT,$ (B2)
where $\varepsilon_{F}$ is generally referred to as the Fermi energy of the system, $a=\pi^{2}/4\varepsilon_{F}$ and $b=\pi^{2}/2\varepsilon_{F}$. Under the low-temperature approximation, $T$ is obtained from Eq. (B1):
$T=\sqrt{\frac{E-E_{\text{min}}}{aN}}.$ (B3)
Using the Boltzmann relation $S=\ln\sigma(E)$ and Eq. (B3), the density of states $\sigma(E)$ can be expressed as:
$\sigma(E)=\text{e}^{S}=\text{e}^{bNT}=\text{e}^{\frac{b}{\sqrt{a}}\sqrt{N(E-E_{\text{min}})}}.$ (B4)
For simplicity, we set $b/\sqrt{a}=1$ when plotting figure 2. However, the qualitative results are unaffected regardless of the value of $b/\sqrt{a}$. The above approach provides an approximation for the density of states of a low-temperature ideal Fermi gas in the limit of $N\to\infty$. For finite $N$, a more detailed discussion of the density of states can be found in Refs. [70, 71].
@@ Appendix C Relative biases of $\beta$ and $T$ in a two-level system
C.1 Relative bias of $\beta$ in a two-level system
For systems with finite $E_{\text{max}}$, Eq. (A2) is modified to
$\begin{split}\langle\hat{\beta}^{(1)}\rangle&=\left.\frac{\sigma(E)\text{e}^{-\beta E}}{Z}\right|^{E_{\text{max}}}_{E_{\text{min}}}-\frac{1}{Z}\int_{E_{\text{min}}}^{E_{\text{max}}}\sigma(E)\text{d}\text{e}^{-\beta E}\\ &=\frac{\sigma(E_{\text{max}})\text{e}^{-\beta E_{\text{max}}}-\sigma(E_{\text{min}})\text{e}^{-\beta E_{\text{min}}}}{Z}+\beta.\end{split}$ (C1)
@@ C.1 Relative bias of $\beta$ in a two-level system
$\sigma(E)=\frac{\Gamma(N+1)}{\varepsilon\Gamma\left(\frac{N}{2}-\frac{E}{\varepsilon}+1\right)\Gamma\left(\frac{N}{2}+\frac{E}{\varepsilon}+1\right)}$ (C2)
and $Z=(\text{e}^{\beta\varepsilon/2}+\text{e}^{-\beta\varepsilon/2})^{N}$. Then, Eq. (C1) is
$\langle\hat{\beta}^{(1)}\rangle=\frac{\text{e}^{\frac{-\beta N\varepsilon}{2}}-\text{e}^{\frac{\beta N\varepsilon}{2}}}{\varepsilon\left(\text{e}^{\frac{\beta\varepsilon}{2}}+\text{e}^{\frac{-\beta\varepsilon}{2}}\right)^{N}}+\beta,$ (C3)
@@
$\begin{split}\Phi_{\beta}&=\left|\frac{\langle\hat{\beta}^{(1)}\rangle-\beta}{\beta}\right|\\ &=\left|\frac{\text{e}^{\frac{-\beta N\varepsilon}{2}}-\text{e}^{\frac{\beta N\varepsilon}{2}}}{\beta\varepsilon\left(\text{e}^{\frac{\beta\varepsilon}{2}}+\text{e}^{\frac{-\beta\varepsilon}{2}}\right)^{N}}\right|\\ &\approx\frac{1}{\beta\varepsilon(1+\text{e}^{-|\beta|\varepsilon})^{N}}\\ &=\frac{1}{\beta\varepsilon}\text{e}^{-N(\ln(1+\text{e}^{-|\beta|\varepsilon}))}\\ &\sim\text{e}^{-\alpha_{1}N},\end{split}$ (C4)
where the approximation used applies $\text{e}^{-N|\beta|\varepsilon}\ll 1$ with large $N$. In figure 6a, $\alpha_{1}$ decreases exponentially with increasing inverse temperature. For a particular $N$, $\Phi_{\beta}$ is non-negligible, which indicates that the unbiased estimator $\hat{\beta}^{(1)}$ remains unbiased except at extremely low temperatures, as given in Eq. (6). This observation is consistent with previous results in quantum thermometry, where the estimation uncertainty diverges exponentially in the low-temperature regime for systems with finite energy gaps [27].
Figure 6: Decay rate varies as NN when $E_{\text{max}}$ is finite. Here, the density of states of an $N$ non-interacting two-level system is Eq. (C2) and we set the spacing of energy levels $\varepsilon=1$.
C.2 Relative bias of $T$ in a two-level system
For systems with finite $E_{\text{max}}$, Eq. (A3) is modified to
$\begin{split}\langle\hat{\beta}^{(-1)}\rangle=&\left.\frac{-\text{d}_{E}^{-1}\sigma(E)\text{e}^{-\beta E}}{Z\beta}\right|^{E_{\text{max}}}_{E_{\text{min}}}+\frac{1}{Z}\int_{E_{\text{min}}}^{E_{\text{max}}}\frac{1}{\beta}\text{e}^{-\beta E}\text{d}\left(\text{d}_{E}^{-1}\sigma(E)\right)\\ =&\frac{\Omega(E_{\text{max}})\text{e}^{-\beta E_{\text{max}}}-\Omega(E_{\text{min}})\text{e}^{-\beta E_{\text{min}}}}{\beta Z}+\beta^{-1}.\end{split}$ (C5)
In an $N$ non-interacting two-level system, we have $\Omega(E_{\text{min}})=\int_{E_{\text{min}}}^{E_{\text{min}}}\sigma(E)\text{d}E=0$ and $\Omega(E_{\text{max}})=\int_{E_{\text{min}}}^{E_{\text{max}}}\sigma(E)\text{d}E=Z|_{\beta=0}$. Then, with the help of Eq. (C2), Eq. (C5) is
$\langle\hat{\beta}^{(-1)}\rangle=\frac{2^{N}\text{e}^{\frac{-\beta N\varepsilon}{2}}}{\beta\left(\text{e}^{\frac{\beta\varepsilon}{2}}+\text{e}^{\frac{-\beta\varepsilon}{2}}\right)^{N}}+\beta^{-1},$ (C6)